A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?
Answer by checkley71(8403)

560/(T+1)=480/T CROSS MULTIPLY
560T=480T+480
560T-480T=480
80T=480
T=480/80
T=6 HOURS IS THE TIME OF THE SECOND LEG OF THE FLIGHT.